If (1/6) – x = 1/6, then x =                                   .
 If (1/6) – x = 1/6, then x =                                   .

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Consider the given equation, (1/6)–x = 1/6
Transpose–x to RHS then it becomes x and 1/6 to LHS then it becomes–1/6
x = 1/6–1/6
x= 0
Then, the value of x = 0

Thanks